OpenStudy (kkbrookly):

Help please. Will give fan and medal (:

OpenStudy (michele_laino):

The velocity v_sof sound within a gas is given by the subsequent formula: \[{v_s} = \sqrt {\frac{{\gamma RT}}{M}} \] where \gamma is a coefficient depending on the gas type, R is the gas constant, T is the absolute temperature, and M is the mass of one mole of that gas.

OpenStudy (michele_laino):

So for a more dense gas, its molecules have lower speed.

OpenStudy (michele_laino):

please wait, I'm searching for your answer, in my textbook

OpenStudy (michele_laino):

we have: \[R = 8.31\quad J/\left( {K \cdot mol} \right)\] is the gas constant

OpenStudy (michele_laino):

whereas \[\gamma >1\] is another constant related to a specific gas. for example for the air we have: \[\gamma = 1.4\]

OpenStudy (michele_laino):

M is the mass of a singl mole of our gas. Namely if N is the number of mole of our gas, and m is the mass of our gas, then we have: \[M = \frac{m}{N}\]

OpenStudy (michele_laino):

finally, T is the absolute temperature of our gas.

OpenStudy (michele_laino):

if t= 0°C, then T = 273.15 K, furthermore, the mass of a molecule of air can be computed if we know the molecular composition of the air, I think that your value is right!

OpenStudy (michele_laino):

oops.. the mass of a mole of air...

OpenStudy (michele_laino):

furthermore, I remember, but I'm not sure, that the coefficient \[\gamma \] is the ratio between the specific heat at constant pressure and the specific heat at constant volume

OpenStudy (michele_laino):

please check my last statement

OpenStudy (michele_laino):

please, wait, the coefficient \[\gamma \], is the same coefficient which is present in the formula of adiabatic transformations, namely: \[p{V^\gamma } = \cos t\] in the pressures-volume plane, so now I'm sure that we have: \[\gamma = \frac{{{c_p}}}{{{c_v}}}\] as I said before.

OpenStudy (michele_laino):

we have to keep in mind that there is an important relationship between C_p and C_v, namely: \[{C_p} - {C_v} = R\]

OpenStudy (michele_laino):

here are some values:

OpenStudy (michele_laino):

monatomic gases like He, Ar \[{C_p} = 5,\quad {C_v} = 3,{\kern 1pt} \quad \gamma = 1.67\]

OpenStudy (michele_laino):

diatomic gases, like H2, N2, O2: \[{C_p} = 7,\quad {C_v} = 5,{\kern 1pt} \quad \gamma = 1.40\]

OpenStudy (michele_laino):

triatomic gases like CO2: \[{C_p} \sim 9,\quad {C_v} \sim 7,{\kern 1pt} \quad \gamma \sim 1.3\]

OpenStudy (michele_laino):

please note that all values are related to one mole of the gas, and their unit of measure is: cal/(mol*K)

OpenStudy (michele_laino):

yes, one value of gamma for He, and another value of gamma, for CO2